Assume you are performing the calculation of (123)
Answer #1

36)

Ones place : 0 in hex (0)

Tens place : 5*2 + 2*1 = 12 in hex (A)

Hundreds place : 5*2 + 5*1 = 15 in hex (F)

Thousands place : 5*1 + 4*1 = 9 in hex (9)

Ten thousand place : 5*1 + 1*0= 5 in hex (5)

Lakhs place : 3*1 + 5*0 = 3 in hex (3)

So hex value is 0x359FA0

34) a=1010 b=0011 so b3 =0

31) c0=a3 = 1

35) a3=1 b3=0

x3=a3 xor b3 = 1 xor 0 = 1

32)

x0= b0 xor a3 = 1 xor 1 = 0

a0 + x0 + c0 = 0 + 0 + 1 now s0 =1 , c1= 0

x1 = b1 xor a3 = 1 xor 1= 0

a1+c1+x1 = 1 + 0 + 1 now s1=0 , c2=1

Hence s1=0

33) a2=0 where a=1010